Anne is 8 years older than Carlos. Betty is 2 years younger that Carlos. The product of their ages is 60. How old is Anne.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Anne is 8 years older than Carlos. Betty is 2 years younger that Carlos. The product of their ages is 60. How old is Anne.
----------
A = C + 8
B = C - 2
A*B*C = 60
============
A, B & C have to all be even, or all be odd.
They can't all be odd, as the product is even --> 3 even integers.
60 = 2*2*3*5
No integer solution.


Answer by ikleyn(52787)

.
The product of WHOSE ages is 60 ?

It is NOT CLEAR from the condition,

and, therefore, this post/(this formulation) IS NOT a Math problem at all.